# Effectively Leveraging Momentum Terms in Stochastic Line Search Frameworks for Fast Optimization of Finite-Sum Problems

### v1

@@ 5 Computational Experiments
We report in Figure 4 and in Figure 5 the comparison of MBCG-DP to its competitors in terms of the training loss and the validation accuracy. The averages of 3 runs with different random seeds are reported. Standard parameters are employed for the monotone line-search of SLS and MSL_SGDM, with $\gamma=0.1$ and $\delta=0.5$, and also for the nonmonotone line search of PoNoS and MBCG-DP, with $\gamma=0.5$ and $\delta=0.5$. For SGD with momentum we use $\alpha=10^{-2},\beta=0.9$ for convex problems and $\alpha=10^{-3},\beta=0.9$ for nonconvex ones. For Adam we use $\alpha=10^{-3}$ for convex problems and $\alpha=10^{-4}$ for nonconvex ones. For both PoNoS and MBCG-DP we use $c=1$ in the definition of the generalized SPS, except for the problem of training the ResNet18 architecture on CIFAR10 dataset, where $c=0.2$ is employed. Reference algorithms utilize a 50% overlap of mini-batches only in configurations specified in Table 2.
@@
As shown in Figure 4, MBCG-DP is extremely effective at reducing the training loss. In all convex tasks it is capable of obtaining high-quality solutions more quickly than the other algorithms considered. In general, the same applies to nonconvex problems, in which MBCG-DP consistently converges faster, except for the problems of training the multi-layer perceptron on MNIST and the ResNet18 on CIFAR10, when the batch size is 128. Looking at trend of the validation accuracy reported in Figure 5, we can note that in these cases MBCG-DP is still capable of reaching top out-of-sample prediction results at least as fast as its competitors. In addition, examining Figure LABEL:fig_acc_cifar10, we can note that the best overall validation accuracy value is reached on the CIFAR10 problem by MBCG-DP with a batch size of 512.
@@ 6 Conclusions
In this work, we introduced an algorithmic framework based on stochastic line searches and momentum directions to tackle nonlinear finite-sum optimization problems in deep learning scenarios. In order to truly benefit from the addition of momentum terms in the search direction, the method leverages the concept of mini-batch persistency. The resulting approach proves to be computationally efficient and effective at solving the class of problems under consideration. In addition, our empirical analysis shows that, for simpler models or small architectures, any algorithm could benefit from the usage of mini-batch persistency.
There are two main research directions that shall be followed at this point: from the one hand, a deep investigation of the convergence properties of algorithms that employ mini-batch persistency would be of significant impact. On the other hand, further computational experiments on even larger models, such as transformer architectures, would be important to assess the usefulness of the proposed approach in the most modern applications.

### v3

@@ 6 Computational Experiments
We report in Figure 5 and in Figure 6 the comparison of MBCG_FR to its competitors in terms of the training loss and the validation accuracy. The algorithms are compared in terms of wall-clock execution time, since employing mini-batch overlap changes the number of updates per epoch, making epoch-based comparison unfair. The averages of 3 runs with different random seeds are reported. Standard parameters are employed for the monotone line-search of SLS and MSL_SGDM, with $\gamma=0.1$ and $\delta=0.5$, and also for the nonmonotone line search of PoNoS and MBCG_FR, with $\gamma=0.5$ and $\delta=0.5$. For SGD with momentum we use $\alpha=10^{-2},\beta=0.9$ for convex problems and $\alpha=10^{-3},\beta=0.9$ for nonconvex ones. For Adam we use $\alpha=10^{-3}$ for convex problems and $\alpha=10^{-4}$ for nonconvex ones. For both PoNoS and MBCG_FR we use $c=1$ in the definition of the generalized SPS, except for the problem of training the ResNet18 architecture on CIFAR10 dataset, where $c=0.2$ is employed. Reference algorithms utilize mini-batch persistency only in configurations specified in Table 2 with the values there reported, whereas MBCG_FR always uses a 50% overlap.
@@
As shown in Figure 5, MBCG_FR is extremely effective at reducing the training loss. In all convex tasks it is capable of obtaining high-quality solutions more quickly than the other algorithms considered. In general, MBCG_FR is also competitive in nonconvex problems, especially for a batch size of 512. Moreover, looking at trend of the validation accuracy reported in Figure 6, we can note that even in cases where Adam is slightly faster, MBCG_FR is capable of reaching top out-of-sample prediction results. In addition, examining Figure LABEL:fig_acc_cifar10, we can note that the best overall validation accuracy value is reached on the CIFAR10 problem by MBCG_FR with a batch size of 512.
@@ 7 Conclusions
In this work, we introduced an algorithmic framework based on stochastic line searches and momentum directions to tackle nonlinear finite-sum optimization problems in deep learning scenarios. In order to truly benefit from the addition of momentum terms in the search direction, the method leverages the concept of mini-batch persistency. The resulting approach, for which we state formal convergence results under suitable assumptions, proves to be computationally efficient and effective at solving the class of problems under consideration. In addition, our empirical analysis shows that, for simpler models or small architectures, any algorithm could benefit from the usage of mini-batch persistency.
There are two main research directions that shall be followed at this point: from the one hand, a deep investigation of the convergence properties of algorithms that employ mini-batch persistency in absence of bias-correction terms would be of significant impact. On the other hand, further computational experiments on even larger models, such as transformer architectures, would be important to assess the usefulness of the proposed approach in the most modern applications.
@@
Acknowledgments
The authors are grateful to the Editor and the reviewers of this manuscript for their constructive comments that helped us to significantly improve the quality of this work. We are very thankful to Professors M. Sciandrone, S. Lucidi and G. Liuzzi for the useful discussions in the early stages of this research.